# ARECHO: Autoregressive Evaluation via Chain-Based Hypothesis Optimization for Speech Multi-Metric Estimation

### v1

@@ Abstract
Speech signal analysis poses significant challenges, particularly in tasks such as speech quality evaluation and profiling, where the goal is to predict multiple perceptual and objective metrics. For instance, metrics like PESQ (Perceptual Evaluation of Speech Quality), STOI (Short-Time Objective Intelligibility), and MOS (Mean Opinion Score) each capture different aspects of speech quality. However, these metrics often have different scales, assumptions, and dependencies, making joint estimation non-trivial. To address these issues, we introduce ARECHO (Autoregressive Evaluation via Chain-based Hypothesis Optimization), a chain-based, versatile evaluation system for speech assessment grounded in autoregressive dependency modeling. ARECHO is distinguished by three key innovations: (1) a comprehensive speech information tokenization pipeline; (2) a dynamic classifier chain that explicitly captures inter-metric dependencies; and (3) a two-step confidence-oriented decoding algorithm that enhances inference reliability. Experiments demonstrate that ARECHO significantly outperforms the baseline framework across diverse evaluation scenarios, including enhanced speech analysis, speech generation evaluation, and, noisy speech evaluation. Furthermore, its dynamic dependency modeling improves interpretability by capturing inter-metric relationships.
@@ 1 Introduction
Challenge I: Diverse Scale Issues. Speech metrics vary significantly in scale and type, complicating joint modeling and optimization. For example, MOS (Mean Opinion Score) ranges from 1 to 5 and SI-SNR is unbounded over (,)($\infty$,$\infty$)( - ∞ , ∞ ). Optimizing such diverse metrics with uniform loss functions (e.g., L1) can lead to biased learning, overemphasizing metrics with larger numerical ranges while under-representing perceptually salient ones like MOS.
@@
To address these challenges, we propose ARECHO (Autoregressive Evaluation via Chain-based Hypothesis Optimization), a flexible and dependency-aware evaluation framework for speech assessment. ARECHO formulates speech evaluation as a chain-based prediction task, utilizing a dynamic classifier chain to model inter-metric relationships in an autoregressive fashion. To handle the diverse scales and types of speech metrics (e.g., categorical vs. continuous, bounded vs. unbounded), we design a speech tokenization pipeline that robustly encodes these heterogeneous metric values into a unified token space. This formulation not only improves predictive performance but also improves interpretability through structured dependency reasoning.
@@ 3.1 Base Task Formulation
$L_{\mathcal{B}}^{i}=L_{\mathcal{B}_{\text{num}}}^{i}+L_{\mathcal{B}_{\text{cat}}}^{i}=\sum_{b\in\mathcal{B}_{\text{num}}}||y_{b}^{i}-\hat{y}_{b}^{i}||_{n}+\sum_{b^{\prime}\in\mathcal{B}_{\text{cat}}}\mathrm{CE}(y_{b^{\prime}}^{i},\hat{y}_{b^{\prime}}^{i}),$ (2)
@@ 3.2 Tokenizing Everything
The full tokenized label sequence for sample $i$ becomes $Z^{i}=\{z^{i}_{b}\}_{b\in\mathcal{B}}$, which serves as the target sequence for autoregressive prediction. For numerical metrics (i.e., $b\in\mathcal{B}_{\text{num}}$), we apply quantization-based tokenization by partitioning the value range into uniformly or adaptively spaced bins. [^4] We experiment with several tokenization strategies including Linear, Normal, Logarithmic, and Quadratic distributions (see Appendix B). While Linear works well overall, Normal and Log-based schemes can be beneficial for metrics with skewed distributions (e.g., SNR, MOS). For categorical metrics (i.e., $b\in\mathcal{B}_{\text{cat}}$), the tokenization is direct by mapping each class label to a unique token.
@@ 3.3 Dynamic Classifier Chain
[…]
@@
where $[b_{1},\dots,b_{K}]$ is a permutation of selected metrics for a given input, and $z^{i}_{b_{k}}$ is the value token for metric $b_{k}$ in sample $i$.
@@
Flexible Inference. At inference time, the model begins with a metadata token (e.g., <Emotion>) and generates the corresponding value token (e.g., Sad). This process repeats for any number of requested metrics, enabling dynamic, chain-style reasoning with arbitrary query orders. This is especially beneficial when metric availability or user interest varies across samples.
@@ 3.4 Two-step Confidence-oriented Decoding
Let $\mathcal{B}_{K}$ be the set of all metrics and $\hat{\mathbf{T}}_{\text{prev}}$ the current decoded prefix after $k$ metrics, corresponding to the already-predicted subset $\mathcal{B}_{k}\subseteq\mathcal{B}_{K}$. With $K-k$ metrics still to predict, the decoding for each remaining metric $b\in\mathcal{B}_{K}\!\setminus\!\mathcal{B}_{k}$ proceeds in two phases:
@@
    and obtain a provisional value token $\hat{z}_{b}$ together with its softmax-based confidence $\text{Conf}(\hat{z}_{b})$.
@@ 4.1 Datasets
Synthesized Speech Data. We utilize two widely recognized benchmark datasets: VoiceMOS 2022 Challenge (Huang et al., 2022) and NISQA (Mittag et al., 2021). VoiceMOS 2022 challenge (Huang et al., 2022) was developed for the VoiceMOS Challenge 2022, comprising a diverse collection of synthetic speech generated by various TTS systems. NISQA (Mittag et al., 2021) dataset also contains synthesized speech samples with corresponding quality ratings across multiple perceptual dimensions.
@@ 4.3 Evaluation
Base Dev. UniVERSA 160.06 0.69 0.27 0.68 0.42
UniVERSA-T 40.95 0.78 0.34 0.70 0.46
ARECHO 25.73 0.86 0.36 0.71 0.51
Enhanced UniVERSA 61.54 0.71 0.27 0.69 0.43
UniVERSA-T 27.34 0.81 0.34 0.70 0.47
ARECHO 20.58 0.84 0.35 0.72 0.51
Corrupted UniVERSA 170.65 0.61 0.24 0.70 0.46
UniVERSA-T 77.72 0.74 0.34 0.71 0.50
ARECHO 44.22 0.82 0.35 0.72 0.55
Synthesized UniVERSA 58.79 0.76 0.27 0.69 0.45
UniVERSA-T 8.10 0.84 0.34 0.72 0.50
ARECHO 4.99 0.91 0.39 0.79 0.65
Avg. Test UniVERSA 96.99 0.69 0.26 0.69 0.45
UniVERSA-T 37.72 0.79 0.34 0.71 0.49
ARECHO 23.26 0.86 0.36 0.74 0.57
Scale Dev. UniVERSA 116.01 0.89 0.37 0.73 0.49
UniVERSA-T 27.98 0.86 0.37 0.74 0.52
ARECHO 29.61 0.85 0.37 0.75 0.52
Enhanced UniVERSA 43.05 0.84 0.34 0.72 0.47
UniVERSA-T 69.94 0.80 0.36 0.74 0.50
ARECHO 32.63 0.83 0.37 0.75 0.53
Corrupted UniVERSA 151.97 0.88 0.37 0.75 0.54
UniVERSA-T 39.80 0.77 0.37 0.76 0.54
ARECHO 34.37 0.84 0.38 0.77 0.56
Synthesized UniVERSA 6.46 0.84 0.32 0.71 0.47
UniVERSA-T 8.23 0.84 0.34 0.73 0.49
ARECHO 8.63 0.85 0.36 0.75 0.54
Avg. Test UniVERSA 67.16 0.85 0.35 0.73 0.50
UniVERSA-T 39.32 0.80 0.36 0.74 0.51
ARECHO 25.21 0.84 0.37 0.76 0.54
@@ 5 Experimental Results
Abalation Analyses. In Appendix I, we present additional studies investigating the effects of tokenization resolution and decoding beam size, further substantiating the flexibility and robustness of the ARECHO framework. [^12] In addition to performance improvements, ARECHO also demonstrates significant gains in training efficiency, which are discussed in detail in Appendix L.
@@
Table 2 highlights how ARECHO internally discovers and exploits an ordering rationale. [^13] While Table 2 presents part of the resulting orders, we further elaborate the details in Appendix J. Across different test sets, metrics related to structured annotations or acoustic scene characteristics (e.g., Q-SpeakerGender, Q-SpeechImpairment, RIR Room Size) consistently appear early in the prediction sequence. These metrics are arguably easier to estimate and provide strong prior information for subsequent metrics. In contrast, more subjective and unstable metrics, such as MOS scores from VoiceMOS and NISQA, tend to appear later in the sequence. This suggests that ARECHO defers harder or noisier predictions until more context is available from previously decoded metrics. Depending on the context, some metrics are also inherently harder to predict e.g., SNR simulation for enhanced and synthesized test sets. For enhanced speech data, the noise was removed by an SE model, while for the synthesized speech, the residual noise in the text-to-speech generated speech signal is usually very low.
@@
As discussed in Sec. 3.3, our proposed dynamic classifier chain supports both flexible order search and inference under arbitrary query sets. In particular, it also allows for inference with a fixed, static order of metrics when such an order is used during training. To investigate how different static orderings affect the performance of ARECHO, we provide additional analyses in the Appendix K.
@@ Appendix B Tokenization with Probability Density Functions
To accommodate heterogeneous evaluation metrics under a unified modeling interface, we transform all metric values into discrete tokens using categorical mapping or quantization-based tokenization. This appendix provides further details on tokenization strategies, especially for numerical metrics.
@@ B.1 Quantization-Based Tokenization for Numerical Metrics
For numerical metrics $b\in\mathcal{B}_{\text{num}}$, we discretize the continuous value space via quantile-based partitioning. Given a collection of observed values $\{y_{b}^{i}\}_{i=1}^{N}$, we define a set of percentiles $\{p_{t}\}_{t=1}^{T-1}$ that divide the empirical distribution into $T$ intervals. The tokenization function $\mathcal{T}_{b}$ maps each $y^{i}_{b}$ into a discrete token $z^{i}_{b}\in\mathcal{V}_{b}$ based on which interval the value belongs to:
$z^{i}_{b}=\mathcal{T}_{b}(y^{i}_{b})=\text{bucket}(y^{i}_{b};\{q_{p_{t}}\}),$ (7)
where $q_{p_{t}}$ denotes the $p_{t}^{\text{th}}$ empirical percentile.
To flexibly control the distribution of token bins, we consider several strategies for generating the percentiles $\{p_{t}\}$:
  • Linear: Uniform spacing across $[0,100]$ percentiles, producing equal-width bins in the cumulative distribution.
  • Exponential: More bins at the high end, emphasizing distinctions among larger values.
  • Logarithmic: More bins at the low end, effective for metrics with sharp variations in small ranges.
  • Normal: Bins concentrated around the mean using a standard Gaussian cumulative distribution.
  • Quadratic: Bins are distributed more heavily at the extremes, capturing boundary effects.
These schemes allow us to modify quantization resolution based on metric characteristics and semantic needs. The resulting token space $\mathcal{V}_{b}$ includes one token per bin, along with reserved tokens for padding and metadata. In our experiments, we use the Linear tokenization.
@@ Appendix C Illustrative Example of Dynamic Classifier Chain
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]

### v2

@@ Abstract
Speech signal analysis poses significant challenges, particularly in tasks such as speech quality evaluation and profiling, where the goal is to predict multiple perceptual and objective metrics. For instance, metrics like PESQ (Perceptual Evaluation of Speech Quality), STOI (Short-Time Objective Intelligibility), and MOS (Mean Opinion Score) each capture different aspects of speech quality. However, these metrics often have different scales, assumptions, and dependencies, making joint estimation non-trivial. To address these issues, we introduce ARECHO (Autoregressive Evaluation via Chain-based Hypothesis Optimization), a chain-based, versatile evaluation system for speech assessment grounded in autoregressive dependency modeling. ARECHO is distinguished by three key innovations: (1) a comprehensive speech information tokenization pipeline; (2) a dynamic classifier chain that explicitly captures inter-metric dependencies; and (3) a two-step confidence-oriented decoding algorithm that enhances inference reliability. Experiments demonstrate that ARECHO significantly outperforms the baseline framework across diverse evaluation scenarios, including enhanced speech analysis, speech generation evaluation, and, noisy speech evaluation. Furthermore, its dynamic dependency modeling improves interpretability by capturing inter-metric relationships. Across tasks, ARECHO offers reference-free evaluation using its dynamic classifier chain to support subset queries (single or multiple metrics) and reduces error propagation via confidence-oriented decoding.
@@ 1 Introduction
Challenge I: Diverse Scale Issues. Speech metrics vary significantly in scale and type, complicating joint modeling and optimization. For example, MOS (Mean Opinion Score) ranges from 1 to 5 and SI-SNR is unbounded over $(-\infty,\infty)$. Optimizing such diverse metrics with uniform loss functions (e.g., L1) can lead to biased learning, overemphasizing metrics with larger numerical ranges while under-representing perceptually salient ones like MOS.
@@
To address these challenges, we propose ARECHO (Autoregressive Evaluation via Chain-based Hypothesis Optimization), a flexible and dependency-aware evaluation framework for speech assessment. ARECHO formulates speech evaluation as a chain-based prediction task, utilizing a dynamic classifier chain to model inter-metric relationships in an autoregressive fashion. To handle the diverse scales and types of speech metrics (e.g., categorical vs. continuous, bounded vs. unbounded), we design a speech tokenization pipeline that robustly encodes these heterogeneous metric values into a unified token space. This formulation not only improves predictive performance but also improves interpretability through structured dependency reasoning. [^1] The code is open-sourced at https://github.com/ftshijt/espnet/tree/universa_plus
@@ 3.1 Base Task Formulation
[…]
@@ 3.2 Tokenizing Everything
The full tokenized label sequence for sample $i$ becomes $Z^{i}=\{z^{i}_{b}\}_{b\in\mathcal{B}}$, which serves as the target sequence for autoregressive prediction. For numerical metrics (i.e., $b\in\mathcal{B}_{\text{num}}$), we apply quantization-based tokenization by partitioning the value range into uniformly or adaptively spaced bins. [^5] Please refer to Appendix B for more discussion.). While Linear works well overall, Normal and Log-based schemes can be beneficial for metrics with skewed distributions (e.g., SNR, MOS). For categorical metrics (i.e., $b\in\mathcal{B}_{\text{cat}}$), the tokenization is direct by mapping each class label to a unique token.
@@ 3.3 Dynamic Classifier Chain
$\mathbf{T}_{\text{full}}=[\texttt{<Gender>},\;\texttt{Male},\;\texttt{<Emotion>},\;\texttt{Happy},\;\texttt{<MOS>},\;\texttt{3.78}]$
@@
where $b_{1},\dots,b_{K}]$ is a permutation of selected metrics for a given input, and $z^{i}_{b_{k}}$ is the value token for metric $b_{k}$ in sample $i$.
@@
Flexible Inference. During inference, the model begins with a metadata token (e.g., <Emotion>) and autoregressively generates its corresponding value token (e.g., Sad). This procedure iterates for any subset of requested metrics, allowing dynamic chain-style reasoning in arbitrary query orders. Users can either (i) specify a fixed prediction order or (ii) rely on our two-step confidence-oriented decoding in Sec. 3.4 to automatically determine an effective sequence. Such flexibility is particularly advantageous when metric availability or user interests differ across samples.
@@ 3.4 Two-step Confidence-oriented Decoding
Let $\mathcal{B}_{K}$ be the set of all metrics and $\hat{\mathbf{T}}_{\text{prev}}$ the current decoded prefix after $k$ metrics, corresponding to the already‐predicted subset $\mathcal{B}_{k}\subseteq\mathcal{B}_{K}$. With KkK-k metrics still to predict, the decoding for each remaining metric $b\in\mathcal{B}_{K}\!\setminus\!\mathcal{B}_{k}$ proceeds in two phases:
@@
    and obtain a provisional value token $\hat{z}_{b}$ together with its softmax‐based confidence $\text{Conf}(\hat{z}_{b})$.
@@ 4.1 Datasets
Synthesized Speech Data. We utilize two benchmark datasets: VoiceMOS 2022 Challenge (Huang et al., 2022) and NISQA (Mittag et al., 2021). VoiceMOS 2022 challenge was developed for the VoiceMOS Challenge 2022, comprising a diverse collection of synthetic speech generated by various TTS systems. NISQA (Mittag et al., 2021) dataset also contains synthesized speech samples with corresponding quality ratings across multiple perceptual dimensions.
@@ 4.3 Evaluation
Base Dev. UniVERSA 160.06 0.69 0.53 0.68 0.42
UniVERSA-T 40.95 0.78 0.68 0.70 0.46
ARECHO 25.73 0.86 0.72 0.71 0.51
Enhanced UniVERSA 61.54 0.71 0.54 0.69 0.43
UniVERSA-T 27.34 0.81 0.68 0.70 0.47
ARECHO 20.58 0.84 0.69 0.72 0.51
Corrupted UniVERSA 170.65 0.61 0.48 0.70 0.46
UniVERSA-T 77.72 0.74 0.67 0.71 0.50
ARECHO 44.22 0.82 0.70 0.72 0.55
Synthesized UniVERSA 58.79 0.76 0.54 0.69 0.45
UniVERSA-T 8.10 0.84 0.68 0.72 0.50
ARECHO 4.99 0.91 0.78 0.79 0.65
Avg. Test UniVERSA 96.99 0.69 0.52 0.69 0.45
UniVERSA-T 37.72 0.79 0.68 0.71 0.49
ARECHO 23.26 0.86 0.72 0.74 0.57
Scale Dev. UniVERSA 116.01 0.89 0.74 0.73 0.49
UniVERSA-T 27.98 0.86 0.75 0.74 0.52
ARECHO 29.61 0.86 0.76 0.75 0.52
Enhanced UniVERSA 43.05 0.84 0.67 0.72 0.47
UniVERSA-T 69.94 0.80 0.71 0.74 0.50
ARECHO 32.63 0.83 0.73 0.75 0.53
Corrupted UniVERSA 151.97 0.88 0.75 0.75 0.54
UniVERSA-T 39.80 0.77 0.74 0.76 0.54
ARECHO 34.37 0.84 0.76 0.77 0.56
Synthesized UniVERSA 6.46 0.84 0.65 0.71 0.47
UniVERSA-T 8.23 0.84 0.68 0.73 0.49
ARECHO 8.63 0.85 0.72 0.75 0.54
Avg. Test UniVERSA 67.16 0.86 0.70 0.73 0.50
UniVERSA-T 39.32 0.82 0.72 0.74 0.51
ARECHO 25.21 0.85 0.74 0.76 0.54
@@ 5 Experimental Results
Ablation and Extended Analysis. We conduct a set of ablations and diagnostic studies to assess robustness, efficiency, and practical behavior; full results are reported in Appendix I and J. These studies examine (i) tokenization resolution, (ii) decoding strategy, (iii) MOS-style perceptual metrics, and (iv) task-oriented metric subsets. Across all settings, ARECHO remains stable: it performs well with compact tokenizations, achieves near-optimal accuracy with simple greedy decoding, and shows strong gains on perceptual quality metrics (e.g., MOS). When trained on either task-specific metric subsets or the full union of metrics, ARECHO shows no evidence of negative transfer from “irrelevant” metrics and often improves on core targets (e.g., SRMR, SDR, human MOS). Together, these results indicate that ARECHO is both effective and adaptive: it can leverage cross-metric structure when helpful while preserving efficiency and specialization. We additionally report efficiency analysis in Appendix M, showing that ARECHO attains these benefits with substantially reduced training and inference cost relative to prior multi-metric systems.
@@
Table 2 highlights how ARECHO internally discovers and exploits an ordering rationale. [^13] While Table 2 presents part of the resulting orders, we further elaborate the details in Appendix K. Across different test sets, metrics related to structured annotations or acoustic scene characteristics (e.g., Q-SpeakerGender, Q-SpeechImpairment, RIR Room Size) consistently appear early in the prediction sequence. These metrics are arguably easier to estimate and provide strong prior information for subsequent metrics. In contrast, more subjective and unstable metrics, such as MOS scores from VoiceMOS and NISQA, tend to appear later in the sequence. This suggests that ARECHO defers harder or noisier predictions until more context is available from previously decoded metrics. Depending on the context, some metrics are also inherently harder to predict e.g., SNR simulation for enhanced and synthesized test sets. For enhanced speech data, the noise was removed by an SE model, while for the synthesized speech, the residual noise in the text-to-speech generated speech signal is usually very low.
@@
As discussed in Sec. 3.3, our proposed dynamic classifier chain supports both flexible order search and inference under arbitrary query sets. In particular, it also allows for inference with a fixed, static order of metrics when such an order is used during training. To investigate how different static orderings affect the performance of ARECHO, we provide additional analyses in the Appendix L.
Complexity and Efficiency. While ARECHO provides several advantages as discussed above, it also maintains modest training cost and strong computational efficiency. By effectively utilizing partially labeled data, ARECHO achieves shorter training time comparable to standard multi-task learning with masking (e.g., UniVERSA). [^14] We present a detailed analysis in Appendix M. Despite its sequential decoding process, ARECHO remains practical: computing the same set of metrics through their original estimators (e.g., LLM-based evaluators) is over $100\times$ slower than a single forward pass of ARECHO.
@@ Appendix B Tokenization with Probability Density Functions
To accommodate heterogeneous evaluation metrics under a unified modeling interface, we transform all metric values into discrete tokens using either categorical mapping or quantization-based tokenization. This appendix elaborates on the numerical-metric tokenization strategies, explains the difference between uniform-over-value and percentile-based binning, and reports a reconstruction study comparing the two.
@@ B.1 Quantization-Based Tokenization for Numerical Metrics
For a numerical metric $b\!\in\!\mathcal{B}_{\text{num}}$, we partition its continuous value domain into $T$ bins, each mapped to a token in the vocabulary $\mathcal{V}_{b}$. Two general binning paradigms are considered:
  • Uniform-over-value binning. The raw range $y_{\min},y_{\max}]$ is divided into $T$ equal-width value intervals. This simple strategy disregards the data distribution and can lead to highly imbalanced token frequencies when the metric values are skewed.
  • Percentile-based binning. Instead of equal value widths, we choose $T$ equally spaced points on the empirical cumulative distribution of the training data. Each bin therefore contains approximately the same number of samples, producing balanced token priors. This strategy (our default) preserves fine detail in dense regions of the data while avoiding sparse bins.
Formally, for percentile-based binning we compute empirical percentiles $\{q_{p_{t}}\}_{t=1}^{T-1}$ with $p_{t}\!=\!100t/T$ and assign each sample value
$z_{b}^{i}=\mathcal{T}_{b}(y_{b}^{i})=\mathrm{bucket}(y_{b}^{i};\{q_{p_{t}}\}),$ (7)
where $\mathrm{bucket}(\cdot)$ returns the index of the interval $\bigl[q_{p_{t-1}},q_{p_{t}}\bigr)$ containing $y_{b}^{i}$. The inverse mapping $\mathcal{T}_{b}^{-1}$ decodes tokens to their bin centroids.
[…]
Domain Tokenization Bins RMSE MAE
Synthesized Percentile 500 0.056 0.014
Uniform 500 0.100 0.025
Percentile 1000 0.034 0.007
Uniform 1000 0.097 0.020
Enhanced Percentile 500 0.139 0.023
Uniform 500 0.223 0.045
Percentile 1000 0.103 0.013
Uniform 1000 0.220 0.039
Corrupted Percentile 500 0.261 0.024
Uniform 500 0.512 0.045
Percentile 1000 0.165 0.012
Uniform 1000 0.507 0.040
According to Table 3, percentile-based tokenization achieves markedly lower reconstruction error than uniform-over-value binning, confirming that balancing token frequencies improves discretization fidelity and generalizes well across domains. Increasing the number of bins from 500 to 1000 further reduces RMSE and MAE, yet we found that such fine granularity makes the autoregressive prediction problem harder (more classes and sparser supervision). Hence, we adopt T=500T{=}500 as a practical trade-off between reconstruction quality and modeling stability.
@@ Appendix C Illustrative Example of Dynamic Classifier Chain
$\texttt{Q-Emotion}=\texttt{"neutral"},\quad\texttt{Q-Background}=\texttt{"indoor"},\quad\texttt{Q-Clarity}=\texttt{"clear"}.$
@@
$\mathbf{T}=[m_{\texttt{Q-Background}},\texttt{"indoor"},m_{\texttt{Q-Clarity}},\texttt{"clear"},m_{\texttt{Q-Emotion}},\texttt{"neutral"}].$
@@
    Input $m_{\texttt{Q-Clarity}}$ → predict $z_{\texttt{Q-Clarity}}=\texttt{"clear"}$
@@
    Input $m_{\texttt{Q-Emotion}}$ → predict $z_{\texttt{Q-Emotion}}=\texttt{"neutral"}$
@@
$\mathbf{T}_{\text{partial}}=[m_{\texttt{Q-Background}},\texttt{"indoor"}].$